Use the Laplace transform to solve the given initial-value problem.

dy/dt  − y = 1,    y(0) = 0

y(t) =



1
Expert's answer
2021-04-14T14:48:29-0400

"\bigstar" Using laplace transform

Theorem - L{f'(t)=sf(s)-f(0)


We have

L("dy\over dt" ) - L (y) = L(1)


sY(s) - y(0)- Y(s)= "1\over s"


(s-1)Y(s)="1\over s"


Y(s)="1\over s(s-1)"

Where Y(s) = L{y(t)}

"\bigstar" In order to find y(t) , we find inverse laplace transform of Y(s) , since


"1\over s(s-1)" = "1\over s-1" - "1\over s"


Now , we have using theorem a= 1 and a=0


y(t) = L-1"1\over s(s-1)" =

=L-1"1\over s-1" - L-1"1\over s"


= et-1

"\boxed{y(t)=e^t-1}"